If (3+2i) + (2+bi) = 5-4i , the value of b is
A)6 B)2 C)-6 D)-2

Answer by amarjeeth123(574)

(3+2i) + (2+bi) = 5-4i
5+(b+2)i=5-4i
Comparing real and imaginary parts we get,
b+2=-4
b=-6
Answer=(C)-6